Connection from IP address 54.196.47.145
Time: 2018 Apr 23 15:20:30
Found 2245 logbook entries, displaying most recent 100

Logbook Contents (Most Recent Entries)

12 Apr 2018 Thursday 16:55 [carter] The beam current meter on the console position 5 has been connected to the EG&G CD1010 current integrator thru a 1 meg ohm resistor. Full scale on the CD1010 will give 1 uA on the console meter position 5. Should connect to the CD1010 monitor output or filter output for a smoother indication.

11 Apr 2018 Wednesday 17:22 [carter] The helium source heater current meter was unstable and has been replaced with a new meter. The old meter had been calibrated to read (heater current / 2) and the new meter will read the actual current in the heater. Temporary calibration currents for the new meter for operating current and standby current are posted on a note located near the meter. When the values are finalized, a new permanent note will be posted.

05 Apr 2018 Thursday 08:10 [Devon] Source cooling water was down to the alarm level. Added water to loop. Alarm light did not clear to off. Tweaked the pot on the sensor power, alarm light cleared.

22 Mar 2018 Thursday 08:24 [Devon] 15L chamber ion gauge developed a leak at one of the filament/grid feedthrus. Replaced.

21 Mar 2018 Wednesday 10:38 [carter/jacobs] Alpha source 24 volt control and TC gauge supply had bad filter cap. Was reading only 20 volts with 4 volts ripple. Capacitor replaced and output voltage restored to 26 volts when in circuit and loaded. Symptom was high reading of filament vacuum on alpha source.

19 Mar 2018 Monday 16:46 [Devon] Alpha source cleaned. Replaced filament vacuum TC gauge.

15 Mar 2018 Thursday 11:03 [carter] Analyzing magnet would not run down. Found Q6 mounting screws (collector connection) had become loose. Tightned and all OK again. May have been cause of intermittend analyzer magnet instability seen once in previous run.

14 Mar 2018 Wednesday 15:40 [carter] In an effort to add stability back to the slit control mode of the TPS, the following adjustment parameters were restored to the factory default values for the slit control system. R6, slit gain modified from 5.0 to 3.1. R12, slit/cpo gain modified from 8.0 to 5.1. R13, slit/cpo crossover frequency modified from 5.0 to 5.8. These adjustments will lower the gain of the slit control system and increase the 3db slit low pass from 8hz to 10hz. Hopefully will restore some stability in slit operation. Some beam/focus combinations were causing oscillations in the control system and required running the control gain at quite low values for stable slit operation.

20 Feb 2018 Tuesday 10:51 [Devon] The current meter/feed wire to chain 3 was laying on the floor. Blue gauge crimp-on ring connector fatigued and broke. Don replaced it with a yellow gauge ring connector, heavier construction. The door is sealed. Gas in Wednesday.

16 Feb 2018 Friday 12:15 [Devon] We lost telemetry from chain 3 during the run today. The terminal dropped some and was unstable. The experiment has been suspended. Gas out Monday.

29 Jan 2018 Monday 11:37 [RSO] accumulated shut down was experienced due to excessive time needed to complete annual calibrations of fixed monitoring system and two hand held neutron detectors

23 Jan 2018 Tuesday 15:41 [Devon] Comfort chiller woes...chiller quit working over winter break, touch screen controls fried preventing operation. FM went ahead and replaced old water pump while chiller was down. Dry loop and cold weather during pump change liberated a great deal of rust in the pipes. Trane replaced the controls but chiller kept shutting down due to poor flow, rust plugged strainer screens, 3.8A on pump motor. Flushed and cleaned screens a number of times, pump flow improved, 4.4A on motor, chiller running. Measured 120-140 gpm with sonic flow meter while air bubbles were circulating. Coils on chiller cleaned. Will continue to monitor motor amps and flush strainers for a few days.

19 Jan 2018 Friday 15:56 [Shiv Subedi] I passed the operator 2 certification with flying colors. :)

12 Jan 2018 Friday 16:14 [Carter] NMR head maintenance. Replaced weak tubes V01 (6922) V02,03 (7788). V04, 05 (5751) OK. Replaced electrolytic capacitors in RF head and stepping motor controller. Also replaced stepping motor because of intermittent lock-up of shaft. Found gears worn and in need of new lubrication. New stepping motor installed. Added correction factors for nmrgoto up and down to restore calibration. Now have very nice NMR signal with original amplitude and symmetry restored. Also lower noise level.

09 Jan 2018 Tuesday 12:21 [Devon] Process chiller loop #1 problem traced back to the pressure switch that controls the fan. Contacts had developed corrosion since it doesn't get used much. Sprayed with LPS lubricant and exorcised the mechanical settings. Fan contactor working smoothly after 2 run cycles.

08 Jan 2018 Monday 10:55 [Devon] I transposed the numbers 3 and 1 in the previous post. #1 is the big circuit, #3 is the small circuit.

07 Jan 2018 Sunday 14:36 [Devon] The process chiller got too cold to run Saturday night. As it warmed up Sunday morning there was some odd behavior, it kept trying to use the big circuit, #3, over and over. I found the #3 pressure driven fan contactor trying to engage and chattering, fan was starting and stopping. I suspect compressor #3 was shutting down on high pressure due to the lack of a fan and never reaching its set point, but, keeping the water below the activation temperature for loops #1 and #2. I've disabled loop #3 and loop #1 has resumed functioning, assume #2 will function if called. Need to investigate the #3 fan contactor and pressure switch this week. And the ACPC CB3 tripped on activation again.

02 Jan 2018 Tuesday 09:44 [Devon] The week before Christmas break we successfully installed new CTI cryopumps and compressor on the HE end of the vault, and replaced the LE tube gate valve.

02 Jan 2018 Tuesday 09:19 [Devon] Pelletron hour meter reading 1/2/2018: 5161. Accelerator operating hours in 2017: 742.

22 Dec 2017 Friday 14:49 [DCI] Pressure between switcher and analyser fallen to 7E-6. If I open to 15degL it falls to 6E-6. Pressure in 15degL rose from 8E-7 to 1.6E-6. If there is a leak from the 45degR it is small.

26 Nov 2017 Sunday 15:24 [Shiv Subedi] Alarm was on for the 'phase failure' in the E1 console and the 'Tritium monitor'. I pressed the 'ack' button in the E1 console and reset button in the tritium monitor. Both of the alarms then went off.

06 Nov 2017 Monday 13:48 [Devon] We are declaring victory on the process chilled water main loo pump. Everything is back to normal with an up-sized motor. A more detailed write up is available in the logbook.

06 Nov 2017 Monday 13:47 [Devon] The 1 hp motor on the process chilled water main loop pump failed spontaneously on a Sunday at 5 pm. Didn’t realize it until Tuesday afternoon when it was noticed the chiller was not running. The alarm threshold was set just below the temperature reached during city water dump cooling. The dump valve opened automatically at a loop temp of 56F and maintained the vacuum systems. The pump bearing assembly had run dry on its oil bath lubrication and the extra drag likely contributed the motor failure. Proceeded to order a replacement 1 hp motor and installed it, also added oil to bearing assembly. The motor was sluggish to start but then ran. Kept an eye on it and found the motor body was 200F after half hour of operation. Shut it down and began to work on replacing the bearing assembly. However, found that the existing bearing assembly and the new one both had basically the same amount of drag when turned by hand. Some, but not much, drag. Put the new motor back on and measured current draw. The new motor was spec’d for 115/230 V, we have 208V. Normal full load at 230V rating is 4.4Amps. We extrapolate that this would be 4.9A at 208V. We measured 5.5A at 208V. Got a transformer to provide 230V and measured 4.6 A at full flow. This motor is working at its 1.15 safety factor limit. The flow in the loop was reduced by partially closing a valve until the current draw by the motor was just below the normal full load reading. In the mean time we used the measured current draw at full loop flow to determine that full flow required 1.3 hp. Ordered a 1.5 hp motor. The bigger motor has been installed and the loop is now flowing at full flow again, motor is drawing 7.5 Amps. 1.3hp/1.5hp = 87%. Rated full load current at 208V is 8.4 A. 7.5A/8.4A = 89%. The full flow load in this loop would appear to be 1.3 hp, too much for the old motor, but not working the new motor too hard.

02 Nov 2017 Thursday 14:00 [Devon] Process chilled water update. A new motor was ordered and installed. The old motor was repaired and works again too. The new motor ran the pump but motor and bearing housing became quite warm after 1/2 hour of operation. Motor is shut off and pump disassembly is underway to replace the bearing assembly.

02 Nov 2017 Thursday 07:23 [DCI] The high water temp/air alarm was sounding this morning and Janitor reported to the police, who called me in. I have disconnected the alarm and reset it. David Ingram

31 Oct 2017 Tuesday 16:53 [Devon] The process chilled water main loop pump quit running sometime this morning. There were no alarms due to...long story. Initial investigation seems like a motor failure. More diagnostics in the morning. We are dumping city water to cool vacuum systems. No magnets, no running until its repaired.

22 Sep 2017 Friday 12:28 [carter] tube lens voltage and tube lens current have been added to the logging webpage plots.

21 Sep 2017 Thursday 16:17 [carter] The settings logger has been upgraded to include the tube lens voltage (tandem.tube_lens_voltage) and tube lens current (tandem.tube_lens_current). The tube lens current is now subtracted from the total LE tube current to indicate only the tandem component off the LE tube current. The currentsum will also indicate correctly again with the Glassman supply. If we return to the Dell supply, the software will need to be modified OR the metering leads on the back of the tube lens control panel (move metering #7 to metering #8, this will log 0 for the lens current).

18 Sep 2017 Monday 09:56 [carter] Any operator that is not NMR certified and would like to be NMR certified should arrange with don for the training. I would hope to complete the training in 1 about hour. Let me know when you are available this week if possible.

17 Sep 2017 Sunday 09:35 [carter] The analyzing magnet power supply has been repaired. Symptoms were that the variac would run down at turn on (normal, run down to lower limit) then activate standby. When supply was turned on the variac would slowly run up some but stop before getting to the 9 volt collector bus pre-regulated voltage. Cause was shorted Q8 causing continuous attempt to run down. Q8 replaced and now supply appears to be functioning normally again. Also, initialy Devon found a blown primary fuse and it was replaced.

12 Sep 2017 Tuesday 10:52 [Tom Massey] We have had problems with the analyzer magnet starting last night. The symptom was that the NMR was not stable. The operator recycled the magnet several times to try and fix the problem. Eventually the NMR was stable enough to continue. This morning the analyzer magnet was found turned down. The magnet was recycled. When it was turned up to the current value for the NMR desired, the magnetic field was much lower than expected. The magnet was recycled again and the current was noted as as maximum of 50 amp rather than 250 amp. The deflection was switched to check for a control problem. The interlock would not come on for either the right or the left deflection. Don Carter was called and he suggested that the problem was due to a blown fuse. Devon checked and replaced the fuses on the analyzer magnet. There was one fuse that was clearly bad. The analyzer magnet now goes to standby and will not interlock. A visual inspection shows several blow resistors. These may be in an abandon part of the circuit or in the current setup.

06 Sep 2017 Wednesday 15:09 [Devon] Athens Fire Department has a training event scheduled here on Thursday, 9/28. Pre planning and confined space rescue.

06 Sep 2017 Wednesday 09:58 [Alexander] Dear operators, we are starting the experiment and need your help with shifts. Please sign up. Thanks !

30 Aug 2017 Wednesday 17:02 [Devon] The tunnel door is repaired.

24 Aug 2017 Thursday 16:27 [meisel] The wireless phones are now back in action and can be used to communicate between the target rooms and the control room.

23 Aug 2017 Wednesday 16:53 [carter] To connect the tube entrance lens to the old Dell power supply follow the following procedure. The Dell supply is located in the pit, near the accelerator tank. 1. Make sure the power to the glassman and Dell supplies are off. The Glassman power switch is located on the front of the supply in the vault. The Dell supply power should be turned off at the circuit breaker on the back of the lense supply control panel in the control room. Leave the Glassman switch OFF when using the Dell supply. In the control room on the back of the tube entrance lens control panel, on the upper terminal strip, replace jumper connections from terminals 1-4, 2-5 and 3-6. This will connect the front panel voltage control pot to the Dell control input. Disconnect the WIRES labeled 4, 5 and 6 from the terminals labeled 4, 5 and 6. This will disconnect the glassman voltage control input to the front panel voltage control pot. Replace the screw-on twinax connectors on J-01, J-02 and J-03 connecting the front panel current and voltage meters from the Dell supply. Disconnect the WIRES labeled 7,8,9 and 10 from terminals 7,8,9 and 10. This will disconnect the glassman current and voltage monitor outputs to the control room panel meters. Now remove the black HV cable from the glassman power supply to the tank and ground the center pin for safety. Connect the large HV cable from the Dell supply to the tank tube lens supply connection where the glassman connection was just removed. Make sure the voltage control in the control room is at zero and turn the circuit breaker on the back of the control room lens supply control panel ON and press the power-on push button. You should be able to control and monitor lens voltage from the control room tube lens voltage control using the Dell supply.

23 Aug 2017 Wednesday 16:51 [carter] To connect the tube entrance lens to the new glassman power supply follow the following procedure. The new glassman 40 KV 7.5 mA supply is located in the small rack enclosure in the vault room across from the pelletron chain drive starter box, near the accelerator tank. The only control on the front panel is an off/on rocker switch that is illuminated when the supply is on. 1. Make sure the power to the glassman and Dell supplies are off. The Glassman power switch is located on the front of the supply in the vault. The Dell supply power should be turned off at the circuit breaker on the back of the lens supply control panel in the control room. Leave this breaker OFF when using the glassman supply. In the control room on the back of the tube entrance lens control panel, on the upper terminal strip, remove jumper connections from terminals 4,5 and 6 but leave in place connected to terminals 1,2 and 3. This will disconnect the front panel voltage control pot to the Dell control input. Connect the WIRES labled 4, 5 and 6 to the terminals labeled 4,5 and 6. This will connect the glassman voltage control input to the front panel voltage control pot. Remove the screw-on twinax connectors on J-01, J-02 and J-03 isolating the front panel current and voltage meters from the Dell supply. Connect the WIRES labeled 7,8,9 and 10 to terminals 7,8,9 and 10. This will connect the glassman current and voltage monitor outputs to the control room panel meters. Now remove the large black HV cable from the Dell power supply to the tank and ground the center pin for safety. Connect the HV cable from the glassman supply to the tank tube lens supply connection where the Dell connection was just removed. Make sure the voltage control in the control room is at zero and turn the power switch on at the glassman supply in the vault. You should be able to control and monitor lens voltage from the control room tube lens voltage control.

23 Aug 2017 Wednesday 16:47 [carter] The Dell tube entrance supply has been repaired and is functioning properly again. It is now available for use but we are currently using the Glassman supply. The Dell supply is kept in reserve in the event of a Glassman supply failure. The instructions for switching between the supplies will follow this log entry. Switching supplies should take about 15 minutes. Currently when using the Glassman supply, the LE tube current will indicate the sum of the lens supply plus the LE tube current. This will upset the current sum monitoring as the real LE tube current will have to subtract the Glassman lens supply current. A future modification to upgrade the logging and monitoring will be made to automatically subtract this lens current from the LE tube current. While using the Glassman supply the power push-button on the lens control panel is not used and power is controlled by a switch on the front of the Glassman supply located in the vault.

23 Aug 2017 Wednesday 15:13 [Dominik] I have completed my operator 2 training!

22 Aug 2017 Tuesday 10:30 [Devon] wifi phone network settings: How to update your Cisco 7925g telephone's WiFi network name Visit https://www.ohio.edu/oit/telephone/faq.cfm for a version of these instructions that includes screen shots of your phone's setup screens. 1. Turn your phone on. 2. Once the phone is on and registered with a phone number, press your phone's Down navigation button to open up Settings. 3. Choose Network Profiles. 4. Select the first profile in your phone's list. It will be named either OU Wireless or Profile 1. 5. Press View. 6. Choose WLAN Configuration. 7. Press View. 8. Wait until your phone displays its WLAN Config screen, then press * * # (in that order) to enable the Change button. 9. Choose SSID and press Change. 10. Replace voip with OHIO University (Make sure to use the exact same capitalization and spelling.) 11. Press Options, then press Save. 12. Choose Security Mode and press Change. 13. Choose PEAP. 14. Press Options, then press Save. 15. Choose UserName and press Change. 16. Enter your OHIO ID. 17. Press Options, then press Save. 18. Choose Password and press Change. 19. Enter your OHIO password. 20. Press Options, then press Save. 21. If you are responsible for more than one phone in your area, you will need to follow the above steps for each one of those phones.

18 Aug 2017 Friday 16:39 [Devon] The wifi cordless telephones in the control room, large target room, vault, and tunnel will likely not work starting Saturday, 8/19. We've been trying to resolve this with IT but were unable to get a solution from IT by the Friday deadline. Will advise when they are working again.

15 Aug 2017 Tuesday 14:48 [Devon] The accelerator and the 15L chamber are back in service. Tank opening notes are available in the logbook.

15 Aug 2017 Tuesday 14:30 [Devon] Tank opening, 4987 hours. Investigation of the negative and unstable upcharge on chain 3. The pulley had shifted sideways on the motor tray due to the bolts on the forward bearing pillow block working loose. There was very little clearance between the chain and the (+) electrode, the chain pellets were striking the electrode. Both the electrode and the pellets suffered chafing. The electrode was sanded clean, NEC suggested sanding and cleaning the chain, but the chafe was not very rough so we left it in place. We will monitor chain performance. The pillow block bolts and holes had grease that had weeped from the bearing above, guessing that helped them work loose. Cleaned as much grease as possible from the holes, installed new bolts with a healthy dose of vibra-tite. The chain 3 motor tray has a static height of 2-7/16" at the alarm micro switch. The switch should activate at around 1-11/16". After a 3/8" drop when running, that leaves 3/8" more before alarm activation. If the chains continue to stretch 3/16" every 2000 hours, we should not see a chain stretch alarm until around 9000 total hours. Chain measurements as follows: Chain 1 static 1-7/8". Previous 2-1/16". 3/16" stretch in 2000 hours, same stretch as the 2000 hours before that. Chain 2 static 1-15/16", previous 2-1/8". Chain 3 static 1-1/16", previous 1-3/16". Chain 1 running 1-9/16". Chain 2 running 1-5/8". Chain 3 running 11/16". The chains have stretch enough that we can remove a link at any time. The chain stretch alarm electronics were found to be fried. The signal had to be moved to E1 panel slot #9. We will add a circuit check to our quarterly test routine.

11 Aug 2017 Friday 16:23 [Devon] The fire alarm was due to a vacuum system malfunction on the 15L chamber. There is a slight smell of burned oil and obviously there was a vapor cloud that was readily visible to the fire alarm optical scanners. The roughing line to the chamber is filled with a clear liquid. The roughing pump still has a full charge of oil. The 15L chamber is out of service until further notice.

10 Aug 2017 Thursday 12:56 [carter] TOF spectrum indicated beam events out of the normal beam burst period, down by about a factor of 1000 but appearing as wrong time by as much as 50 ns. Re-tuning chopping and bunching controls had little if no effect. Went back to DC tune and found low energy X (2X or inflection magnet) was not peaked. Tuning inflection magnet with 2X set at 5.0 gained about 30% more beam. Chopped and bunched after DC tune and found the out of time blob was not present anymore.

04 Aug 2017 Friday 16:49 [carter/ingram] A test was made of the chain charging current monitoring system by injecting a known current (100 uA) at the tank feedthru for each of the three chains. All chain current meters read a stable 100 uA so it appears that the chain current monitoring system outside the tank is functioning properly.

04 Aug 2017 Friday 13:07 [Devon] There is a malfunction in the accelerator charging system. The charging controller is reporting negative charge on chain 3 and the total upcharge is erratic. However the terminal voltage seemed unaffected. Did a quick check with chain 3 running alone, sounded like it was running, and the terminal rose when suppressor/inductor voltage was applied, although the terminal power supplies were not in play since the generator was not being driven. We have switched to running with chains 1 & 2 only in hopes of finishing the C12 run. The charging readings are stable but the suppressor/inductor voltages are higher to compensate for the lack of chain 3. This configuration limits our maximum voltage to a TBD value. Further evaluation will be needed to see if we need a tank entry, or if this is an external metering problem.

26 Jul 2017 Wednesday 09:35 [devon] The sputter source developed a dead short on the target plane last night. I tried the reaming probe thru the target air lock but that didn't clear the short. This morning I vented the source and removed the target plane with argon flowing. Found: 1) precipitate plating on one of the ionizer straps had broken loose on one end and curled inwards in a springy fashion to touch the outside of the cathode sheath; 2) the insulator was clean and unstressed. Precipitate was vacuumed out of the ionizer cavity and source sealed up. When the source recovers to high vacuum it will be activated and a new D pellet installed since the current one is near the bottom. Watch the status page for beam production...

25 Jul 2017 Tuesday 16:31 [Tom Massey] Cody Parker, Doug Soltez and Shiv Subedi have completed "Tritium Awareness "Training.

21 Jul 2017 Friday 12:31 [Devon] The massive tunnel door has suffered a major failure and is inoperable. Investigation and repair planning is underway, but this will not be a very quick fix. Accelerator and source shut down.

20 Jul 2017 Thursday 15:31 [Justin Warren] I completed Operator 2 training!

17 Jul 2017 Monday 15:31 [Doug Soltesz] Wanted to let everybody know that I completed my operator 2 training. WOOOO!

11 Jul 2017 Tuesday 12:03 [Devon] Replaced the ECO O2 sensors on the VESDA system. Next set due for replacement by 6/20/2019. Order from Orr Protection, not Simplex. Simplex order was at 12 weeks and they couldn't provide any status. Orr order was 3 1/2 weeks.

06 Jul 2017 Thursday 19:51 [Tom Massey] We have completed the measruements for Daren Bleuel from LLNL. Thanks to everyone who has helped in this work. Especial thanks to Don Carter and Devon Jacobs for perservering through all of the problems.

06 Jul 2017 Thursday 16:39 [Tom Massey] I wish to note that we started with a 33 microampere beam of deuterons in the LE cup this morining. This is very goood source performance.

05 Jul 2017 Wednesday 18:27 [Tom Massey] Around 6:15, the sputter source tripped off. The Emergency off button was red, the Dell High voltage and the Glassman high voltages were off. No alarms were sounding. There was lightning from the thunderstorm. A possible cause. Target and Extractor buttons pushed to drop the voltage on startup. Ionizer turned to 10 A. Cs Resivoir and Cs Feed turned off as temperature dropped.

05 Jul 2017 Wednesday 08:09 [Devon] SF6 alarm was sounding this morning. Did a quick leak check before starting the source, tank joint is giving a good hit. Mechanical gauge says we are down 1 psi since last gas transfer. HVAC trouble may be contributing too.

04 Jul 2017 Tuesday 09:59 [Devon] The chiller was factory repaired Monday evening and is functioning normally. Lab systems (except for the 15L chamber) are back to normal and the run can resume as soon as beam comes up Wednesday. The 15L chamber seems to have a pretty good leak somewhere.

03 Jul 2017 Monday 12:59 [Devon] Chiller: Factory service did not show up as advertised today. Unclear why, or when they will, but Wednesday perhaps? Many people on holiday today. The chiller has been running at 50% today, but no guarantee it will stay that way until it is repaired. We continue to keep proccess sytems in a warm water state...

02 Jul 2017 Sunday 09:07 [Devon] The chiller prediction was correct, it shut down. All process chilled pipes dripping with water. Raised all the sub loop temperature settings to get above dew point. Turned off magnets. Analyzer cryo pump seems to have crashed. Valved off and turned off 15L diff pumps since they may not have liked the higher water temps.

01 Jul 2017 Saturday 13:55 [Tom Massey] The emergency crew has done what they can and have called TRANE to be here on Monday. They are unable to clear completely two sensors to allow the unit to run. One is the high pressure cuttoff and the other was the High Pressure in the Coolant Inlet. They have checked the coolant pressures and have made up for any lost coolant. They were unable to connect to the controlling program to reset the system. The chiller has been restarted but is not expected to run long. We will not try to run until Monday.

01 Jul 2017 Saturday 12:39 [Tom Massey] Facilies management has sent three people to work in the Comfort Chiller. The HVAC person in Jeff. The codes read out as high intake pressure. The screens were cleaned, an inline water filter was taken out and cleaned. Make up water was added. They are working on the controls. They are having trouble keeping the chiller running under controls.

01 Jul 2017 Saturday 10:27 [Tom Massey] The comfort chiller has stopped working. The chiller is now warming up. The facilities management emergency number was called and we are waitng for someone to come.

30 Jun 2017 Friday 20:44 [Devon] Darren B. came back from dinner to find the source low flow alarm active. The cause was that the source water pump turned off for some reason, which takes out the source power platform. Nothing else in the building went out. There are reports of a sizeable AEP outage in the area. The pump restatred and continues to run. The source is back on to continue running in the new pellet.

30 Jun 2017 Friday 18:31 [Devon] We changed the source pellet and beam production stalled. Shut down the accelerator and closed gate valves. Letting source run overnight. Beam meter is on 10 uA scale...

29 Jun 2017 Thursday 21:55 [Andrea] The Emergency Off button near the Swinger was pressed while attepting to align the targets in the chamber. This was unintentional and the alarm condition has been reset.

28 Jun 2017 Wednesday 15:58 [Devon] The beam in the swinger became unstable and eventually unusable last evening. The solder link alarm was activating when the swinger cup was operated. Inspection of the swinger cup this morning found that the current lead wire had broken by fatigue. The cup was charging and discharging since it was floating electrically. To compound the issues, the lead wire had a fiberglass insulating jacket on it, and the loose wire and jacket just happened to come to rest on the beam axis blocking beam. Situation has been corrected.

23 Jun 2017 Friday 13:47 [Devon] The beam line is repaired and the leak check was non-detect. We had a bellows section that used to be on the swinger when it had a cryopump. The tubes will need to be monitored and conditioned while going up in voltage next week.

22 Jun 2017 Thursday 15:53 [Devon] There is a vacuum leak in a bellows section of beam pipe at the HE end of the accelerator. The accelerator controls have been disabled and the cryopumps valved off. The leak will vent the tubes to atm overnight for repairs tomorrow.

22 Jun 2017 Thursday 14:47 [Devon] Sputter source cesium reservoir ran dry, bone dry. Glass shards were crystal clear. Feed tube clean. Replaced Cs. Last load was Sept 2016.

15 Jun 2017 Thursday 15:43 [Devon] The police now have access to the restricted area door at the bottom of the main stairs. The police, or the fire department, will only enter the restricted area in certain emergency situations. Otherwise they will wait in the lobby for an operator or keyholder as before. You may see the Standard Operating Procedure (SOP) for when they can and can't enter in the Handbook, Emergency Procedures section. Please read it. There is also a new Emergency Off button by the drinking fountain. It is for them to use if they need to come in and the accelerator is still on with no operator available. You can of course use this button on your way out if something is forcing you to flee the control room.

14 Jun 2017 Wednesday 11:16 [devon] It appears that 120 C is the maximum temperature for producing a carbon beam with the sputter source. Tried 125 and extractor current surged too high.

07 Jun 2017 Wednesday 15:57 [carter] Test message sent to "all operators list" including new operators.

04 Jun 2017 Sunday 06:37 [carter] Modified Tank Volume SF6 hi-res plot on real-time web page to display 32-36 instead of 33-37 bottles SF6.

16 May 2017 Tuesday 13:45 [devon] this is an example

04 May 2017 Thursday 15:20 [carter] The H.E Slit current meter on the Terminal Potential Stabilizer has been replaced. It should not stick on zero anymore. If you notice problems let me know.

29 Apr 2017 Saturday 16:36 [Ingram] Source water was off, so the aplha source power was off. Source water pump restarted. Power restored to the alpha source. All looks good.

29 Apr 2017 Saturday 16:30 [Ingram] There was a power outage this morning that tripped a bunch of alarms that Kristyn reset. Many things did not appear to be affected. But, the pump valve on the HE extension was closed. I reopened it. The valve between the analyzer and switcher magnets was also closed. This valve took a long time to open/operate, about a minute. Also, one filament on the ion gauge is blown. I hope we have a spare, because the next filament failure will require the gauge to be changed.

28 Apr 2017 Friday 11:26 [carter/brandenburg] A concrete density measurement was made by weighing (68.4 Kg, 151 lb.) round concrete core drill slug that came from the control room ceiling. The dimensions were 15.75 in. diameter by 9.5 in. high resulting in a calculated concrete density of 2.33 +- .04 g/cm3. Architectural drawings listed all structural concrete used in the lab construction was 4500 psi compression where concrete used on grade and footers was 3500 psi. This should be the 4500 psi compression concrete.

24 Apr 2017 Monday 15:42 [Devon] The sputter source burned across its extractor insulator. Source has been cleaned. Insulator replaced. Fresh Boron pellet loaded. The existing cesium was preserved under Argon. The "outgassing" procedure will need to be done prior to starting the source. The TV camera is pointed at the source ion gauge vacuum. It can take a long time to get beam from a clean source, don't be afraid to let it run overnight.

17 Apr 2017 Monday 16:15 [Devon] Discovered the cause of the weekend fire alarm in the LTR. The 65R leg gate valve solenoid fried.

16 Apr 2017 Sunday 00:38 [Devon] Fire alarm in LTR pipe 3 (middle bay) and elevated detections on other pipes and control room. Major stink of burned electrical. Emergency lighting CFL bulb in middle of LTR was out, could be culprit. Ran exhaust fans, stink gone, ready for easter bunny.

14 Apr 2017 Friday 11:19 [Alexander] Dear operators, we are going to resume the 11B+48Ca experiment on Monday, April 17, so help with shifts is greatly appreciated. Thanks !

12 Apr 2017 Wednesday 15:53 [carter/jacobs] The HVEC corona current feedthru started leaking SF6 and was also sparking at the designated grid voltage of -10 V. At a grid voltage of -10 volts, the 6BK4C corona current regulator tube plate was 17.5 KV and the corona feedthru was arcing to ground. The original feedthru seemed very fragile and it was decided to replace it with a coaxial cable type compression feedthru that had also been supplied by HVEC for another tank high voltage feedthru and was modified to accept RG-217/U cable and should have a DC breakdown in excess of 50KV. The internal spring/wire spool mechanism was modified to accept the solid core wire of the RG-217/U cable into a receiving socket so that the cable can be replaced without removing the corona point assembly. The short length of cable is inserted into the compression fitting, slides into the center conductor receiving socket and protrudes about 4" outside the tank and terminates into a 5/8" ball spark gap spaced from ground 1.1 cm for a spark gap voltage of approx. 40 KV. A new length of Glassman RG-8/U cable connects at the feedthru spark gap ball and runs to the corona current control module that contains the 6BK4C control tube. There is a spark gap at the plate of the 6BK4C tube inside the NEC corona control unit that consists of two 1/2" brass spheres separated by 1" giving an estimated spark gap breakdown of 80KV. The 6BK4C specifies a max plate voltage of 27 KV. We operate at a nominal corona current of 35uA. The original target grid voltage was -10 volts. At -10 volts on the grid of the 6BK4C the plate voltage was 19.5 KV. At times of instability the control grid could reach -25 volts and the 6BK4C plate voltage would reach 40KV and spark at the original feedthru. A modification was made to the corona control unit to limit the grid voltage to a maximum grid voltage of -15 volts and the 6BK4C plate would be limited to 24.5 KV to protect the tube and the feedthru. Since the stability of the pelletron is very smooth we have modified the standard operating point of the grid voltage to be -5 volts instead of -10 volts. This will lower the nominal voltage at the feedthru from 17.5 KV to 9.2 KV. If more dynamic range of control is needed in the future, the -10V operating point may be safely used at 17.5 KV.

06 Apr 2017 Thursday 13:02 [Devon] The active shooter training (ALiCE) will be offered on both Tuesdays, 4/11 & 4/18, from 9:00-12:00. The first 2 hours will be the general ALiCE training that could help you at anytime and anyplace in this day and age. The third hour will be to discuss the unique hazards that could arise in the accelerator lab. I have to admit, during the planning meeting for this training, a lot of good, useful, suggestions came up. This is not a mandatory training and its not just for operators and keyholders, tell your associates in the building too.

04 Apr 2017 Tuesday 13:04 [Devon] As requested, trying to set up active shooter training for everyone. Possible dates are Tuesday, 4/11, 9:00-12:00, or Tuesday, 4/18, same time. More info to follow...

31 Mar 2017 Friday 11:09 [carter] Dr. Andreas Ruben of W-IE-NE-R Plein & Baus Corp will be in the R.W Finlay Conference Room at the Edwards Accelerator Lab on Monday 3 April 2017 at 10:00 AM to demonstrate new equipment offerings they have for nuclear data acquisition and control needs. See http://www.wiener-us.com for info on their products. Please spread the word to anyone you feel might benefit from the demonstration.

28 Mar 2017 Tuesday 14:08 [Devon] The high voltage feedthru on the corona probe is failing. As a precaution the SF6 in the accelerator is being taken down to 0 psig (1 atm) to prevent gas loss. This means that we are extremely limited in capabilities until we are ready to make a permanent repair. It is our hope that we can still support the lab class for RBS, but there will be special operational parameters. You can read more about the saga in the logbook.

28 Mar 2017 Tuesday 14:02 [Devon] There was a spark after 2 hours of operation. The grid was running at -5 which corresponds to 10 KV at the corona feedthrough. Normal operation would give 18 KV at the feedthrough. We measured current and voltage on the corona feedthrough and found the values to be as expected, however the feedthrough continued to spark as soon as we reached 15 KV. Then we discovered that the feedthrough is also leaking SF6. The leak pegged both ranges on the new detector and still registered on the less sensitive old detector. Taking the gas down to 1 atm to prevent loss by catastrophic failure.

27 Mar 2017 Monday 15:45 [Devon] We’ve had a spell of spark trouble. Trying to run at 4 MV and repeated sparks. Reached the point that sparks were happening at 3.4 MV. But no sign in the data of resistor noise, charging system working fine, all indications that everything is normal. Began testing. Retracted the corona points and ran quite nicely at 4.1 MV. Moved the corona points in to operating distance started controlling – spark. Tried again, spark without controlling. Moved points back out, everything fine. Put a grounded current meter directly on the corona feedthrough to eliminate tube and HV cable, everything fine while passing 35 uA of current. Found the Teflon (?) insulator that was around the feedthrough had a spark hole in it. Shortened the Teflon insulator so that burn spot was not at the HV point, drilled out the burned spot too. Put everything back and spark. Changed the HV cable and tube, spark again. Cleaned the feedthrough and removed the Teflon (?) insulator and it all seems to be working at 4.017 MV. Experiment resuming…

22 Mar 2017 Wednesday 18:29 [Alexander] Dear operators We are starting a new experiment with the 11B 20 MeV beam on 48Ca target. This is collaborative experimet with faculty and students from Central Michigan University. They will come on Friday, March 24 and we will be running from Friday, March 24 to Friday March 31. I would like to ask students to help with shifts (including weekend March 25-26, if possible) and also encourage students to come and talk to visitors and see what is going on. Thank you.

16 Feb 2017 Thursday 09:15 [carter] Surveyed small target room and could locate no failed equipment. Tested DAQ systems and module electronics and all seem OK. Campus closed for water main break.

16 Feb 2017 Thursday 03:30 [Devon] There was a fire alarm that originated in the small target room at 02:28. There was also minor detections in Vault pipe 1. FD and PD were waiting when I arrived. On inspection, there was a distinct smell of burned electrical compents in the STR but smoke levels were back to normal. Unable to identify the source of the burn. Mech and Diff pumps still working, lights all working. Need to be on the lookout for for failed electrical component....

02 Feb 2017 Thursday 13:01 [Devon] I have scheduled fire alarm testing for Monday, 2/6, at 8:00 am.

30 Jan 2017 Monday 11:02 [Alexander] Fisrt of all, many thanks to Kristyn Brandenburg, Andrea Richard and Rekam Giri for their great help with shifts last week, and would like to remind operators that we are running this week as well and still need to fill in 3 evening shifts on Tuesday, Wednesday and Friday (01/31, 02/01, and 02/03/2017). Thank you !

27 Jan 2017 Friday 17:16 [carter] A new Epson PowerLite 955WH LCD Projector has been installed in the R.W.Finlay conference room. It is brighter, better resolution and has more inputs (HDMI also). Please give it a try before attempting to give a presentation and get familiar with it. The new HDMI cable is labeled "overhead HDMI 1". Other input cables are the same as before.

20 Jan 2017 Friday 10:51 [carter,massey,brooks] Radiation monitor yearly calibration check, each detector at 10, 20, 40 mR OK. 80% alarm check OK and 100% machine shutdown OK.

20 Jan 2017 Friday 10:09 [Alexander] We are going to start a new run on Monday, January 23, with a deuteron 0.5 MeV beam on Tritium target. This will be a continuation of the experiment on the deuteron break up cross section measurements in support of experiments at the National Ignition Facility of Lawrence Livermore Lab. We will try to keep it running for two weeks, so help from operators is needed. You are welcome to join the experiment by signing up for shifts , help on evening shifts is more preferable but other time slots are also welcomed. Thanks !

19 Jan 2017 Thursday 15:34 [Devon] Tank Opening, December 2016. --- 4419 chain hours. --- Searching for source of instability in GVM, tube/column downcharges, and tube entrance lens. Suspicious of gas quality and recirculation effectiveness. --- Inspected LE Tube grid entrance lens, nothing remarkable. Cleaned both LE Penning gauges. --- New blower to recirculate SF6. Old blower had much wear, would have taken more shop time than the cost of a new one. Also new pressure relief valve for blower, set for 6 psi. Old PRV was worn and chattered when running, filled pig with brass flakes. Source of noisy recirculator for how many years? New blower is producing 4 psi head, therefore PRV not opening in normal use. Very quiet. New chain coupling also. --- Corona points replaced. Current set from last entry had worn points with mild deposits. Replaced with tungsten points. 0.13" protrusion. 0.0125 - 0.1563 is range. --- Added selector valves and a pressure transducer to each of the 3 filter canisters in the recirculation loop. Replaced filters. Tried to seal filter cans with 4401 thermoseal gasket material but small leaks were present. Seal surfaces not smoothly machined. Switched to viton gaskets. Pressure drop data as follows: Blower in/out: 4.1 psi, Tank in/out: 1.2 psi, Dryer/Scrubber: 1.9psi, V1 Filter: 0.3 psi, V19 Filter: 0.3 psi, Dryer Filter: 0.3 psi. --- Changed Scrubber sieve material. All 3 sieves gained weight. 5A 11%, 4A 6%, 3A 9%. --- Inspected all resistors, all good. --- In the end it turned out that the GVM grounding brush was the problem, see separate notes on that.

19 Jan 2017 Thursday 13:36 [carter] Radiation monitor vault detector zero unstable. This amplifier was the last remaining original Kiethley K302 electrometer amplifier. It has now been replaced with a Burr Brown OPA128 electrometer amplifier as the others have been replaced.

Connection from IP address 54.196.47.145
Time: 2018 Apr 23 15:20:30
Found 2245 logbook entries, displaying most recent 100

Logbook Contents (Most Recent Entries)

12 Apr 2018 Thursday 16:55 [carter] The beam current meter on the console position 5 has been connected to the EG&G CD1010 current integrator thru a 1 meg ohm resistor. Full scale on the CD1010 will give 1 uA on the console meter position 5. Should connect to the CD1010 monitor output or filter output for a smoother indication.

11 Apr 2018 Wednesday 17:22 [carter] The helium source heater current meter was unstable and has been replaced with a new meter. The old meter had been calibrated to read (heater current / 2) and the new meter will read the actual current in the heater. Temporary calibration currents for the new meter for operating current and standby current are posted on a note located near the meter. When the values are finalized, a new permanent note will be posted.

05 Apr 2018 Thursday 08:10 [Devon] Source cooling water was down to the alarm level. Added water to loop. Alarm light did not clear to off. Tweaked the pot on the sensor power, alarm light cleared.

22 Mar 2018 Thursday 08:24 [Devon] 15L chamber ion gauge developed a leak at one of the filament/grid feedthrus. Replaced.

21 Mar 2018 Wednesday 10:38 [carter/jacobs] Alpha source 24 volt control and TC gauge supply had bad filter cap. Was reading only 20 volts with 4 volts ripple. Capacitor replaced and output voltage restored to 26 volts when in circuit and loaded. Symptom was high reading of filament vacuum on alpha source.

19 Mar 2018 Monday 16:46 [Devon] Alpha source cleaned. Replaced filament vacuum TC gauge.

15 Mar 2018 Thursday 11:03 [carter] Analyzing magnet would not run down. Found Q6 mounting screws (collector connection) had become loose. Tightned and all OK again. May have been cause of intermittend analyzer magnet instability seen once in previous run.

14 Mar 2018 Wednesday 15:40 [carter] In an effort to add stability back to the slit control mode of the TPS, the following adjustment parameters were restored to the factory default values for the slit control system. R6, slit gain modified from 5.0 to 3.1. R12, slit/cpo gain modified from 8.0 to 5.1. R13, slit/cpo crossover frequency modified from 5.0 to 5.8. These adjustments will lower the gain of the slit control system and increase the 3db slit low pass from 8hz to 10hz. Hopefully will restore some stability in slit operation. Some beam/focus combinations were causing oscillations in the control system and required running the control gain at quite low values for stable slit operation.

20 Feb 2018 Tuesday 10:51 [Devon] The current meter/feed wire to chain 3 was laying on the floor. Blue gauge crimp-on ring connector fatigued and broke. Don replaced it with a yellow gauge ring connector, heavier construction. The door is sealed. Gas in Wednesday.

16 Feb 2018 Friday 12:15 [Devon] We lost telemetry from chain 3 during the run today. The terminal dropped some and was unstable. The experiment has been suspended. Gas out Monday.

29 Jan 2018 Monday 11:37 [RSO] accumulated shut down was experienced due to excessive time needed to complete annual calibrations of fixed monitoring system and two hand held neutron detectors

23 Jan 2018 Tuesday 15:41 [Devon] Comfort chiller woes...chiller quit working over winter break, touch screen controls fried preventing operation. FM went ahead and replaced old water pump while chiller was down. Dry loop and cold weather during pump change liberated a great deal of rust in the pipes. Trane replaced the controls but chiller kept shutting down due to poor flow, rust plugged strainer screens, 3.8A on pump motor. Flushed and cleaned screens a number of times, pump flow improved, 4.4A on motor, chiller running. Measured 120-140 gpm with sonic flow meter while air bubbles were circulating. Coils on chiller cleaned. Will continue to monitor motor amps and flush strainers for a few days.

19 Jan 2018 Friday 15:56 [Shiv Subedi] I passed the operator 2 certification with flying colors. :)

12 Jan 2018 Friday 16:14 [Carter] NMR head maintenance. Replaced weak tubes V01 (6922) V02,03 (7788). V04, 05 (5751) OK. Replaced electrolytic capacitors in RF head and stepping motor controller. Also replaced stepping motor because of intermittent lock-up of shaft. Found gears worn and in need of new lubrication. New stepping motor installed. Added correction factors for nmrgoto up and down to restore calibration. Now have very nice NMR signal with original amplitude and symmetry restored. Also lower noise level.

09 Jan 2018 Tuesday 12:21 [Devon] Process chiller loop #1 problem traced back to the pressure switch that controls the fan. Contacts had developed corrosion since it doesn't get used much. Sprayed with LPS lubricant and exorcised the mechanical settings. Fan contactor working smoothly after 2 run cycles.

08 Jan 2018 Monday 10:55 [Devon] I transposed the numbers 3 and 1 in the previous post. #1 is the big circuit, #3 is the small circuit.

07 Jan 2018 Sunday 14:36 [Devon] The process chiller got too cold to run Saturday night. As it warmed up Sunday morning there was some odd behavior, it kept trying to use the big circuit, #3, over and over. I found the #3 pressure driven fan contactor trying to engage and chattering, fan was starting and stopping. I suspect compressor #3 was shutting down on high pressure due to the lack of a fan and never reaching its set point, but, keeping the water below the activation temperature for loops #1 and #2. I've disabled loop #3 and loop #1 has resumed functioning, assume #2 will function if called. Need to investigate the #3 fan contactor and pressure switch this week. And the ACPC CB3 tripped on activation again.

02 Jan 2018 Tuesday 09:44 [Devon] The week before Christmas break we successfully installed new CTI cryopumps and compressor on the HE end of the vault, and replaced the LE tube gate valve.

02 Jan 2018 Tuesday 09:19 [Devon] Pelletron hour meter reading 1/2/2018: 5161. Accelerator operating hours in 2017: 742.

22 Dec 2017 Friday 14:49 [DCI] Pressure between switcher and analyser fallen to 7E-6. If I open to 15degL it falls to 6E-6. Pressure in 15degL rose from 8E-7 to 1.6E-6. If there is a leak from the 45degR it is small.

26 Nov 2017 Sunday 15:24 [Shiv Subedi] Alarm was on for the 'phase failure' in the E1 console and the 'Tritium monitor'. I pressed the 'ack' button in the E1 console and reset button in the tritium monitor. Both of the alarms then went off.

06 Nov 2017 Monday 13:48 [Devon] We are declaring victory on the process chilled water main loo pump. Everything is back to normal with an up-sized motor. A more detailed write up is available in the logbook.

06 Nov 2017 Monday 13:47 [Devon] The 1 hp motor on the process chilled water main loop pump failed spontaneously on a Sunday at 5 pm. Didn’t realize it until Tuesday afternoon when it was noticed the chiller was not running. The alarm threshold was set just below the temperature reached during city water dump cooling. The dump valve opened automatically at a loop temp of 56F and maintained the vacuum systems. The pump bearing assembly had run dry on its oil bath lubrication and the extra drag likely contributed the motor failure. Proceeded to order a replacement 1 hp motor and installed it, also added oil to bearing assembly. The motor was sluggish to start but then ran. Kept an eye on it and found the motor body was 200F after half hour of operation. Shut it down and began to work on replacing the bearing assembly. However, found that the existing bearing assembly and the new one both had basically the same amount of drag when turned by hand. Some, but not much, drag. Put the new motor back on and measured current draw. The new motor was spec’d for 115/230 V, we have 208V. Normal full load at 230V rating is 4.4Amps. We extrapolate that this would be 4.9A at 208V. We measured 5.5A at 208V. Got a transformer to provide 230V and measured 4.6 A at full flow. This motor is working at its 1.15 safety factor limit. The flow in the loop was reduced by partially closing a valve until the current draw by the motor was just below the normal full load reading. In the mean time we used the measured current draw at full loop flow to determine that full flow required 1.3 hp. Ordered a 1.5 hp motor. The bigger motor has been installed and the loop is now flowing at full flow again, motor is drawing 7.5 Amps. 1.3hp/1.5hp = 87%. Rated full load current at 208V is 8.4 A. 7.5A/8.4A = 89%. The full flow load in this loop would appear to be 1.3 hp, too much for the old motor, but not working the new motor too hard.

02 Nov 2017 Thursday 14:00 [Devon] Process chilled water update. A new motor was ordered and installed. The old motor was repaired and works again too. The new motor ran the pump but motor and bearing housing became quite warm after 1/2 hour of operation. Motor is shut off and pump disassembly is underway to replace the bearing assembly.

02 Nov 2017 Thursday 07:23 [DCI] The high water temp/air alarm was sounding this morning and Janitor reported to the police, who called me in. I have disconnected the alarm and reset it. David Ingram

31 Oct 2017 Tuesday 16:53 [Devon] The process chilled water main loop pump quit running sometime this morning. There were no alarms due to...long story. Initial investigation seems like a motor failure. More diagnostics in the morning. We are dumping city water to cool vacuum systems. No magnets, no running until its repaired.

22 Sep 2017 Friday 12:28 [carter] tube lens voltage and tube lens current have been added to the logging webpage plots.

21 Sep 2017 Thursday 16:17 [carter] The settings logger has been upgraded to include the tube lens voltage (tandem.tube_lens_voltage) and tube lens current (tandem.tube_lens_current). The tube lens current is now subtracted from the total LE tube current to indicate only the tandem component off the LE tube current. The currentsum will also indicate correctly again with the Glassman supply. If we return to the Dell supply, the software will need to be modified OR the metering leads on the back of the tube lens control panel (move metering #7 to metering #8, this will log 0 for the lens current).

18 Sep 2017 Monday 09:56 [carter] Any operator that is not NMR certified and would like to be NMR certified should arrange with don for the training. I would hope to complete the training in 1 about hour. Let me know when you are available this week if possible.

17 Sep 2017 Sunday 09:35 [carter] The analyzing magnet power supply has been repaired. Symptoms were that the variac would run down at turn on (normal, run down to lower limit) then activate standby. When supply was turned on the variac would slowly run up some but stop before getting to the 9 volt collector bus pre-regulated voltage. Cause was shorted Q8 causing continuous attempt to run down. Q8 replaced and now supply appears to be functioning normally again. Also, initialy Devon found a blown primary fuse and it was replaced.

12 Sep 2017 Tuesday 10:52 [Tom Massey] We have had problems with the analyzer magnet starting last night. The symptom was that the NMR was not stable. The operator recycled the magnet several times to try and fix the problem. Eventually the NMR was stable enough to continue. This morning the analyzer magnet was found turned down. The magnet was recycled. When it was turned up to the current value for the NMR desired, the magnetic field was much lower than expected. The magnet was recycled again and the current was noted as as maximum of 50 amp rather than 250 amp. The deflection was switched to check for a control problem. The interlock would not come on for either the right or the left deflection. Don Carter was called and he suggested that the problem was due to a blown fuse. Devon checked and replaced the fuses on the analyzer magnet. There was one fuse that was clearly bad. The analyzer magnet now goes to standby and will not interlock. A visual inspection shows several blow resistors. These may be in an abandon part of the circuit or in the current setup.

06 Sep 2017 Wednesday 15:09 [Devon] Athens Fire Department has a training event scheduled here on Thursday, 9/28. Pre planning and confined space rescue.

06 Sep 2017 Wednesday 09:58 [Alexander] Dear operators, we are starting the experiment and need your help with shifts. Please sign up. Thanks !

30 Aug 2017 Wednesday 17:02 [Devon] The tunnel door is repaired.

24 Aug 2017 Thursday 16:27 [meisel] The wireless phones are now back in action and can be used to communicate between the target rooms and the control room.

23 Aug 2017 Wednesday 16:53 [carter] To connect the tube entrance lens to the old Dell power supply follow the following procedure. The Dell supply is located in the pit, near the accelerator tank. 1. Make sure the power to the glassman and Dell supplies are off. The Glassman power switch is located on the front of the supply in the vault. The Dell supply power should be turned off at the circuit breaker on the back of the lense supply control panel in the control room. Leave the Glassman switch OFF when using the Dell supply. In the control room on the back of the tube entrance lens control panel, on the upper terminal strip, replace jumper connections from terminals 1-4, 2-5 and 3-6. This will connect the front panel voltage control pot to the Dell control input. Disconnect the WIRES labeled 4, 5 and 6 from the terminals labeled 4, 5 and 6. This will disconnect the glassman voltage control input to the front panel voltage control pot. Replace the screw-on twinax connectors on J-01, J-02 and J-03 connecting the front panel current and voltage meters from the Dell supply. Disconnect the WIRES labeled 7,8,9 and 10 from terminals 7,8,9 and 10. This will disconnect the glassman current and voltage monitor outputs to the control room panel meters. Now remove the black HV cable from the glassman power supply to the tank and ground the center pin for safety. Connect the large HV cable from the Dell supply to the tank tube lens supply connection where the glassman connection was just removed. Make sure the voltage control in the control room is at zero and turn the circuit breaker on the back of the control room lens supply control panel ON and press the power-on push button. You should be able to control and monitor lens voltage from the control room tube lens voltage control using the Dell supply.

23 Aug 2017 Wednesday 16:51 [carter] To connect the tube entrance lens to the new glassman power supply follow the following procedure. The new glassman 40 KV 7.5 mA supply is located in the small rack enclosure in the vault room across from the pelletron chain drive starter box, near the accelerator tank. The only control on the front panel is an off/on rocker switch that is illuminated when the supply is on. 1. Make sure the power to the glassman and Dell supplies are off. The Glassman power switch is located on the front of the supply in the vault. The Dell supply power should be turned off at the circuit breaker on the back of the lens supply control panel in the control room. Leave this breaker OFF when using the glassman supply. In the control room on the back of the tube entrance lens control panel, on the upper terminal strip, remove jumper connections from terminals 4,5 and 6 but leave in place connected to terminals 1,2 and 3. This will disconnect the front panel voltage control pot to the Dell control input. Connect the WIRES labled 4, 5 and 6 to the terminals labeled 4,5 and 6. This will connect the glassman voltage control input to the front panel voltage control pot. Remove the screw-on twinax connectors on J-01, J-02 and J-03 isolating the front panel current and voltage meters from the Dell supply. Connect the WIRES labeled 7,8,9 and 10 to terminals 7,8,9 and 10. This will connect the glassman current and voltage monitor outputs to the control room panel meters. Now remove the large black HV cable from the Dell power supply to the tank and ground the center pin for safety. Connect the HV cable from the glassman supply to the tank tube lens supply connection where the Dell connection was just removed. Make sure the voltage control in the control room is at zero and turn the power switch on at the glassman supply in the vault. You should be able to control and monitor lens voltage from the control room tube lens voltage control.

23 Aug 2017 Wednesday 16:47 [carter] The Dell tube entrance supply has been repaired and is functioning properly again. It is now available for use but we are currently using the Glassman supply. The Dell supply is kept in reserve in the event of a Glassman supply failure. The instructions for switching between the supplies will follow this log entry. Switching supplies should take about 15 minutes. Currently when using the Glassman supply, the LE tube current will indicate the sum of the lens supply plus the LE tube current. This will upset the current sum monitoring as the real LE tube current will have to subtract the Glassman lens supply current. A future modification to upgrade the logging and monitoring will be made to automatically subtract this lens current from the LE tube current. While using the Glassman supply the power push-button on the lens control panel is not used and power is controlled by a switch on the front of the Glassman supply located in the vault.

23 Aug 2017 Wednesday 15:13 [Dominik] I have completed my operator 2 training!

22 Aug 2017 Tuesday 10:30 [Devon] wifi phone network settings: How to update your Cisco 7925g telephone's WiFi network name Visit https://www.ohio.edu/oit/telephone/faq.cfm for a version of these instructions that includes screen shots of your phone's setup screens. 1. Turn your phone on. 2. Once the phone is on and registered with a phone number, press your phone's Down navigation button to open up Settings. 3. Choose Network Profiles. 4. Select the first profile in your phone's list. It will be named either OU Wireless or Profile 1. 5. Press View. 6. Choose WLAN Configuration. 7. Press View. 8. Wait until your phone displays its WLAN Config screen, then press * * # (in that order) to enable the Change button. 9. Choose SSID and press Change. 10. Replace voip with OHIO University (Make sure to use the exact same capitalization and spelling.) 11. Press Options, then press Save. 12. Choose Security Mode and press Change. 13. Choose PEAP. 14. Press Options, then press Save. 15. Choose UserName and press Change. 16. Enter your OHIO ID. 17. Press Options, then press Save. 18. Choose Password and press Change. 19. Enter your OHIO password. 20. Press Options, then press Save. 21. If you are responsible for more than one phone in your area, you will need to follow the above steps for each one of those phones.

18 Aug 2017 Friday 16:39 [Devon] The wifi cordless telephones in the control room, large target room, vault, and tunnel will likely not work starting Saturday, 8/19. We've been trying to resolve this with IT but were unable to get a solution from IT by the Friday deadline. Will advise when they are working again.

15 Aug 2017 Tuesday 14:48 [Devon] The accelerator and the 15L chamber are back in service. Tank opening notes are available in the logbook.

15 Aug 2017 Tuesday 14:30 [Devon] Tank opening, 4987 hours. Investigation of the negative and unstable upcharge on chain 3. The pulley had shifted sideways on the motor tray due to the bolts on the forward bearing pillow block working loose. There was very little clearance between the chain and the (+) electrode, the chain pellets were striking the electrode. Both the electrode and the pellets suffered chafing. The electrode was sanded clean, NEC suggested sanding and cleaning the chain, but the chafe was not very rough so we left it in place. We will monitor chain performance. The pillow block bolts and holes had grease that had weeped from the bearing above, guessing that helped them work loose. Cleaned as much grease as possible from the holes, installed new bolts with a healthy dose of vibra-tite. The chain 3 motor tray has a static height of 2-7/16" at the alarm micro switch. The switch should activate at around 1-11/16". After a 3/8" drop when running, that leaves 3/8" more before alarm activation. If the chains continue to stretch 3/16" every 2000 hours, we should not see a chain stretch alarm until around 9000 total hours. Chain measurements as follows: Chain 1 static 1-7/8". Previous 2-1/16". 3/16" stretch in 2000 hours, same stretch as the 2000 hours before that. Chain 2 static 1-15/16", previous 2-1/8". Chain 3 static 1-1/16", previous 1-3/16". Chain 1 running 1-9/16". Chain 2 running 1-5/8". Chain 3 running 11/16". The chains have stretch enough that we can remove a link at any time. The chain stretch alarm electronics were found to be fried. The signal had to be moved to E1 panel slot #9. We will add a circuit check to our quarterly test routine.

11 Aug 2017 Friday 16:23 [Devon] The fire alarm was due to a vacuum system malfunction on the 15L chamber. There is a slight smell of burned oil and obviously there was a vapor cloud that was readily visible to the fire alarm optical scanners. The roughing line to the chamber is filled with a clear liquid. The roughing pump still has a full charge of oil. The 15L chamber is out of service until further notice.

10 Aug 2017 Thursday 12:56 [carter] TOF spectrum indicated beam events out of the normal beam burst period, down by about a factor of 1000 but appearing as wrong time by as much as 50 ns. Re-tuning chopping and bunching controls had little if no effect. Went back to DC tune and found low energy X (2X or inflection magnet) was not peaked. Tuning inflection magnet with 2X set at 5.0 gained about 30% more beam. Chopped and bunched after DC tune and found the out of time blob was not present anymore.

04 Aug 2017 Friday 16:49 [carter/ingram] A test was made of the chain charging current monitoring system by injecting a known current (100 uA) at the tank feedthru for each of the three chains. All chain current meters read a stable 100 uA so it appears that the chain current monitoring system outside the tank is functioning properly.

04 Aug 2017 Friday 13:07 [Devon] There is a malfunction in the accelerator charging system. The charging controller is reporting negative charge on chain 3 and the total upcharge is erratic. However the terminal voltage seemed unaffected. Did a quick check with chain 3 running alone, sounded like it was running, and the terminal rose when suppressor/inductor voltage was applied, although the terminal power supplies were not in play since the generator was not being driven. We have switched to running with chains 1 & 2 only in hopes of finishing the C12 run. The charging readings are stable but the suppressor/inductor voltages are higher to compensate for the lack of chain 3. This configuration limits our maximum voltage to a TBD value. Further evaluation will be needed to see if we need a tank entry, or if this is an external metering problem.

26 Jul 2017 Wednesday 09:35 [devon] The sputter source developed a dead short on the target plane last night. I tried the reaming probe thru the target air lock but that didn't clear the short. This morning I vented the source and removed the target plane with argon flowing. Found: 1) precipitate plating on one of the ionizer straps had broken loose on one end and curled inwards in a springy fashion to touch the outside of the cathode sheath; 2) the insulator was clean and unstressed. Precipitate was vacuumed out of the ionizer cavity and source sealed up. When the source recovers to high vacuum it will be activated and a new D pellet installed since the current one is near the bottom. Watch the status page for beam production...

25 Jul 2017 Tuesday 16:31 [Tom Massey] Cody Parker, Doug Soltez and Shiv Subedi have completed "Tritium Awareness "Training.

21 Jul 2017 Friday 12:31 [Devon] The massive tunnel door has suffered a major failure and is inoperable. Investigation and repair planning is underway, but this will not be a very quick fix. Accelerator and source shut down.

20 Jul 2017 Thursday 15:31 [Justin Warren] I completed Operator 2 training!

17 Jul 2017 Monday 15:31 [Doug Soltesz] Wanted to let everybody know that I completed my operator 2 training. WOOOO!

11 Jul 2017 Tuesday 12:03 [Devon] Replaced the ECO O2 sensors on the VESDA system. Next set due for replacement by 6/20/2019. Order from Orr Protection, not Simplex. Simplex order was at 12 weeks and they couldn't provide any status. Orr order was 3 1/2 weeks.

06 Jul 2017 Thursday 19:51 [Tom Massey] We have completed the measruements for Daren Bleuel from LLNL. Thanks to everyone who has helped in this work. Especial thanks to Don Carter and Devon Jacobs for perservering through all of the problems.

06 Jul 2017 Thursday 16:39 [Tom Massey] I wish to note that we started with a 33 microampere beam of deuterons in the LE cup this morining. This is very goood source performance.

05 Jul 2017 Wednesday 18:27 [Tom Massey] Around 6:15, the sputter source tripped off. The Emergency off button was red, the Dell High voltage and the Glassman high voltages were off. No alarms were sounding. There was lightning from the thunderstorm. A possible cause. Target and Extractor buttons pushed to drop the voltage on startup. Ionizer turned to 10 A. Cs Resivoir and Cs Feed turned off as temperature dropped.

05 Jul 2017 Wednesday 08:09 [Devon] SF6 alarm was sounding this morning. Did a quick leak check before starting the source, tank joint is giving a good hit. Mechanical gauge says we are down 1 psi since last gas transfer. HVAC trouble may be contributing too.

04 Jul 2017 Tuesday 09:59 [Devon] The chiller was factory repaired Monday evening and is functioning normally. Lab systems (except for the 15L chamber) are back to normal and the run can resume as soon as beam comes up Wednesday. The 15L chamber seems to have a pretty good leak somewhere.

03 Jul 2017 Monday 12:59 [Devon] Chiller: Factory service did not show up as advertised today. Unclear why, or when they will, but Wednesday perhaps? Many people on holiday today. The chiller has been running at 50% today, but no guarantee it will stay that way until it is repaired. We continue to keep proccess sytems in a warm water state...

02 Jul 2017 Sunday 09:07 [Devon] The chiller prediction was correct, it shut down. All process chilled pipes dripping with water. Raised all the sub loop temperature settings to get above dew point. Turned off magnets. Analyzer cryo pump seems to have crashed. Valved off and turned off 15L diff pumps since they may not have liked the higher water temps.

01 Jul 2017 Saturday 13:55 [Tom Massey] The emergency crew has done what they can and have called TRANE to be here on Monday. They are unable to clear completely two sensors to allow the unit to run. One is the high pressure cuttoff and the other was the High Pressure in the Coolant Inlet. They have checked the coolant pressures and have made up for any lost coolant. They were unable to connect to the controlling program to reset the system. The chiller has been restarted but is not expected to run long. We will not try to run until Monday.

01 Jul 2017 Saturday 12:39 [Tom Massey] Facilies management has sent three people to work in the Comfort Chiller. The HVAC person in Jeff. The codes read out as high intake pressure. The screens were cleaned, an inline water filter was taken out and cleaned. Make up water was added. They are working on the controls. They are having trouble keeping the chiller running under controls.

01 Jul 2017 Saturday 10:27 [Tom Massey] The comfort chiller has stopped working. The chiller is now warming up. The facilities management emergency number was called and we are waitng for someone to come.

30 Jun 2017 Friday 20:44 [Devon] Darren B. came back from dinner to find the source low flow alarm active. The cause was that the source water pump turned off for some reason, which takes out the source power platform. Nothing else in the building went out. There are reports of a sizeable AEP outage in the area. The pump restatred and continues to run. The source is back on to continue running in the new pellet.

30 Jun 2017 Friday 18:31 [Devon] We changed the source pellet and beam production stalled. Shut down the accelerator and closed gate valves. Letting source run overnight. Beam meter is on 10 uA scale...

29 Jun 2017 Thursday 21:55 [Andrea] The Emergency Off button near the Swinger was pressed while attepting to align the targets in the chamber. This was unintentional and the alarm condition has been reset.

28 Jun 2017 Wednesday 15:58 [Devon] The beam in the swinger became unstable and eventually unusable last evening. The solder link alarm was activating when the swinger cup was operated. Inspection of the swinger cup this morning found that the current lead wire had broken by fatigue. The cup was charging and discharging since it was floating electrically. To compound the issues, the lead wire had a fiberglass insulating jacket on it, and the loose wire and jacket just happened to come to rest on the beam axis blocking beam. Situation has been corrected.

23 Jun 2017 Friday 13:47 [Devon] The beam line is repaired and the leak check was non-detect. We had a bellows section that used to be on the swinger when it had a cryopump. The tubes will need to be monitored and conditioned while going up in voltage next week.

22 Jun 2017 Thursday 15:53 [Devon] There is a vacuum leak in a bellows section of beam pipe at the HE end of the accelerator. The accelerator controls have been disabled and the cryopumps valved off. The leak will vent the tubes to atm overnight for repairs tomorrow.

22 Jun 2017 Thursday 14:47 [Devon] Sputter source cesium reservoir ran dry, bone dry. Glass shards were crystal clear. Feed tube clean. Replaced Cs. Last load was Sept 2016.

15 Jun 2017 Thursday 15:43 [Devon] The police now have access to the restricted area door at the bottom of the main stairs. The police, or the fire department, will only enter the restricted area in certain emergency situations. Otherwise they will wait in the lobby for an operator or keyholder as before. You may see the Standard Operating Procedure (SOP) for when they can and can't enter in the Handbook, Emergency Procedures section. Please read it. There is also a new Emergency Off button by the drinking fountain. It is for them to use if they need to come in and the accelerator is still on with no operator available. You can of course use this button on your way out if something is forcing you to flee the control room.

14 Jun 2017 Wednesday 11:16 [devon] It appears that 120 C is the maximum temperature for producing a carbon beam with the sputter source. Tried 125 and extractor current surged too high.

07 Jun 2017 Wednesday 15:57 [carter] Test message sent to "all operators list" including new operators.

04 Jun 2017 Sunday 06:37 [carter] Modified Tank Volume SF6 hi-res plot on real-time web page to display 32-36 instead of 33-37 bottles SF6.

16 May 2017 Tuesday 13:45 [devon] this is an example

04 May 2017 Thursday 15:20 [carter] The H.E Slit current meter on the Terminal Potential Stabilizer has been replaced. It should not stick on zero anymore. If you notice problems let me know.

29 Apr 2017 Saturday 16:36 [Ingram] Source water was off, so the aplha source power was off. Source water pump restarted. Power restored to the alpha source. All looks good.

29 Apr 2017 Saturday 16:30 [Ingram] There was a power outage this morning that tripped a bunch of alarms that Kristyn reset. Many things did not appear to be affected. But, the pump valve on the HE extension was closed. I reopened it. The valve between the analyzer and switcher magnets was also closed. This valve took a long time to open/operate, about a minute. Also, one filament on the ion gauge is blown. I hope we have a spare, because the next filament failure will require the gauge to be changed.

28 Apr 2017 Friday 11:26 [carter/brandenburg] A concrete density measurement was made by weighing (68.4 Kg, 151 lb.) round concrete core drill slug that came from the control room ceiling. The dimensions were 15.75 in. diameter by 9.5 in. high resulting in a calculated concrete density of 2.33 +- .04 g/cm3. Architectural drawings listed all structural concrete used in the lab construction was 4500 psi compression where concrete used on grade and footers was 3500 psi. This should be the 4500 psi compression concrete.

24 Apr 2017 Monday 15:42 [Devon] The sputter source burned across its extractor insulator. Source has been cleaned. Insulator replaced. Fresh Boron pellet loaded. The existing cesium was preserved under Argon. The "outgassing" procedure will need to be done prior to starting the source. The TV camera is pointed at the source ion gauge vacuum. It can take a long time to get beam from a clean source, don't be afraid to let it run overnight.

17 Apr 2017 Monday 16:15 [Devon] Discovered the cause of the weekend fire alarm in the LTR. The 65R leg gate valve solenoid fried.

16 Apr 2017 Sunday 00:38 [Devon] Fire alarm in LTR pipe 3 (middle bay) and elevated detections on other pipes and control room. Major stink of burned electrical. Emergency lighting CFL bulb in middle of LTR was out, could be culprit. Ran exhaust fans, stink gone, ready for easter bunny.

14 Apr 2017 Friday 11:19 [Alexander] Dear operators, we are going to resume the 11B+48Ca experiment on Monday, April 17, so help with shifts is greatly appreciated. Thanks !

12 Apr 2017 Wednesday 15:53 [carter/jacobs] The HVEC corona current feedthru started leaking SF6 and was also sparking at the designated grid voltage of -10 V. At a grid voltage of -10 volts, the 6BK4C corona current regulator tube plate was 17.5 KV and the corona feedthru was arcing to ground. The original feedthru seemed very fragile and it was decided to replace it with a coaxial cable type compression feedthru that had also been supplied by HVEC for another tank high voltage feedthru and was modified to accept RG-217/U cable and should have a DC breakdown in excess of 50KV. The internal spring/wire spool mechanism was modified to accept the solid core wire of the RG-217/U cable into a receiving socket so that the cable can be replaced without removing the corona point assembly. The short length of cable is inserted into the compression fitting, slides into the center conductor receiving socket and protrudes about 4" outside the tank and terminates into a 5/8" ball spark gap spaced from ground 1.1 cm for a spark gap voltage of approx. 40 KV. A new length of Glassman RG-8/U cable connects at the feedthru spark gap ball and runs to the corona current control module that contains the 6BK4C control tube. There is a spark gap at the plate of the 6BK4C tube inside the NEC corona control unit that consists of two 1/2" brass spheres separated by 1" giving an estimated spark gap breakdown of 80KV. The 6BK4C specifies a max plate voltage of 27 KV. We operate at a nominal corona current of 35uA. The original target grid voltage was -10 volts. At -10 volts on the grid of the 6BK4C the plate voltage was 19.5 KV. At times of instability the control grid could reach -25 volts and the 6BK4C plate voltage would reach 40KV and spark at the original feedthru. A modification was made to the corona control unit to limit the grid voltage to a maximum grid voltage of -15 volts and the 6BK4C plate would be limited to 24.5 KV to protect the tube and the feedthru. Since the stability of the pelletron is very smooth we have modified the standard operating point of the grid voltage to be -5 volts instead of -10 volts. This will lower the nominal voltage at the feedthru from 17.5 KV to 9.2 KV. If more dynamic range of control is needed in the future, the -10V operating point may be safely used at 17.5 KV.

06 Apr 2017 Thursday 13:02 [Devon] The active shooter training (ALiCE) will be offered on both Tuesdays, 4/11 & 4/18, from 9:00-12:00. The first 2 hours will be the general ALiCE training that could help you at anytime and anyplace in this day and age. The third hour will be to discuss the unique hazards that could arise in the accelerator lab. I have to admit, during the planning meeting for this training, a lot of good, useful, suggestions came up. This is not a mandatory training and its not just for operators and keyholders, tell your associates in the building too.

04 Apr 2017 Tuesday 13:04 [Devon] As requested, trying to set up active shooter training for everyone. Possible dates are Tuesday, 4/11, 9:00-12:00, or Tuesday, 4/18, same time. More info to follow...

31 Mar 2017 Friday 11:09 [carter] Dr. Andreas Ruben of W-IE-NE-R Plein & Baus Corp will be in the R.W Finlay Conference Room at the Edwards Accelerator Lab on Monday 3 April 2017 at 10:00 AM to demonstrate new equipment offerings they have for nuclear data acquisition and control needs. See http://www.wiener-us.com for info on their products. Please spread the word to anyone you feel might benefit from the demonstration.

28 Mar 2017 Tuesday 14:08 [Devon] The high voltage feedthru on the corona probe is failing. As a precaution the SF6 in the accelerator is being taken down to 0 psig (1 atm) to prevent gas loss. This means that we are extremely limited in capabilities until we are ready to make a permanent repair. It is our hope that we can still support the lab class for RBS, but there will be special operational parameters. You can read more about the saga in the logbook.

28 Mar 2017 Tuesday 14:02 [Devon] There was a spark after 2 hours of operation. The grid was running at -5 which corresponds to 10 KV at the corona feedthrough. Normal operation would give 18 KV at the feedthrough. We measured current and voltage on the corona feedthrough and found the values to be as expected, however the feedthrough continued to spark as soon as we reached 15 KV. Then we discovered that the feedthrough is also leaking SF6. The leak pegged both ranges on the new detector and still registered on the less sensitive old detector. Taking the gas down to 1 atm to prevent loss by catastrophic failure.

27 Mar 2017 Monday 15:45 [Devon] We’ve had a spell of spark trouble. Trying to run at 4 MV and repeated sparks. Reached the point that sparks were happening at 3.4 MV. But no sign in the data of resistor noise, charging system working fine, all indications that everything is normal. Began testing. Retracted the corona points and ran quite nicely at 4.1 MV. Moved the corona points in to operating distance started controlling – spark. Tried again, spark without controlling. Moved points back out, everything fine. Put a grounded current meter directly on the corona feedthrough to eliminate tube and HV cable, everything fine while passing 35 uA of current. Found the Teflon (?) insulator that was around the feedthrough had a spark hole in it. Shortened the Teflon insulator so that burn spot was not at the HV point, drilled out the burned spot too. Put everything back and spark. Changed the HV cable and tube, spark again. Cleaned the feedthrough and removed the Teflon (?) insulator and it all seems to be working at 4.017 MV. Experiment resuming…

22 Mar 2017 Wednesday 18:29 [Alexander] Dear operators We are starting a new experiment with the 11B 20 MeV beam on 48Ca target. This is collaborative experimet with faculty and students from Central Michigan University. They will come on Friday, March 24 and we will be running from Friday, March 24 to Friday March 31. I would like to ask students to help with shifts (including weekend March 25-26, if possible) and also encourage students to come and talk to visitors and see what is going on. Thank you.

16 Feb 2017 Thursday 09:15 [carter] Surveyed small target room and could locate no failed equipment. Tested DAQ systems and module electronics and all seem OK. Campus closed for water main break.

16 Feb 2017 Thursday 03:30 [Devon] There was a fire alarm that originated in the small target room at 02:28. There was also minor detections in Vault pipe 1. FD and PD were waiting when I arrived. On inspection, there was a distinct smell of burned electrical compents in the STR but smoke levels were back to normal. Unable to identify the source of the burn. Mech and Diff pumps still working, lights all working. Need to be on the lookout for for failed electrical component....

02 Feb 2017 Thursday 13:01 [Devon] I have scheduled fire alarm testing for Monday, 2/6, at 8:00 am.

30 Jan 2017 Monday 11:02 [Alexander] Fisrt of all, many thanks to Kristyn Brandenburg, Andrea Richard and Rekam Giri for their great help with shifts last week, and would like to remind operators that we are running this week as well and still need to fill in 3 evening shifts on Tuesday, Wednesday and Friday (01/31, 02/01, and 02/03/2017). Thank you !

27 Jan 2017 Friday 17:16 [carter] A new Epson PowerLite 955WH LCD Projector has been installed in the R.W.Finlay conference room. It is brighter, better resolution and has more inputs (HDMI also). Please give it a try before attempting to give a presentation and get familiar with it. The new HDMI cable is labeled "overhead HDMI 1". Other input cables are the same as before.

20 Jan 2017 Friday 10:51 [carter,massey,brooks] Radiation monitor yearly calibration check, each detector at 10, 20, 40 mR OK. 80% alarm check OK and 100% machine shutdown OK.

20 Jan 2017 Friday 10:09 [Alexander] We are going to start a new run on Monday, January 23, with a deuteron 0.5 MeV beam on Tritium target. This will be a continuation of the experiment on the deuteron break up cross section measurements in support of experiments at the National Ignition Facility of Lawrence Livermore Lab. We will try to keep it running for two weeks, so help from operators is needed. You are welcome to join the experiment by signing up for shifts , help on evening shifts is more preferable but other time slots are also welcomed. Thanks !

19 Jan 2017 Thursday 15:34 [Devon] Tank Opening, December 2016. --- 4419 chain hours. --- Searching for source of instability in GVM, tube/column downcharges, and tube entrance lens. Suspicious of gas quality and recirculation effectiveness. --- Inspected LE Tube grid entrance lens, nothing remarkable. Cleaned both LE Penning gauges. --- New blower to recirculate SF6. Old blower had much wear, would have taken more shop time than the cost of a new one. Also new pressure relief valve for blower, set for 6 psi. Old PRV was worn and chattered when running, filled pig with brass flakes. Source of noisy recirculator for how many years? New blower is producing 4 psi head, therefore PRV not opening in normal use. Very quiet. New chain coupling also. --- Corona points replaced. Current set from last entry had worn points with mild deposits. Replaced with tungsten points. 0.13" protrusion. 0.0125 - 0.1563 is range. --- Added selector valves and a pressure transducer to each of the 3 filter canisters in the recirculation loop. Replaced filters. Tried to seal filter cans with 4401 thermoseal gasket material but small leaks were present. Seal surfaces not smoothly machined. Switched to viton gaskets. Pressure drop data as follows: Blower in/out: 4.1 psi, Tank in/out: 1.2 psi, Dryer/Scrubber: 1.9psi, V1 Filter: 0.3 psi, V19 Filter: 0.3 psi, Dryer Filter: 0.3 psi. --- Changed Scrubber sieve material. All 3 sieves gained weight. 5A 11%, 4A 6%, 3A 9%. --- Inspected all resistors, all good. --- In the end it turned out that the GVM grounding brush was the problem, see separate notes on that.

19 Jan 2017 Thursday 13:36 [carter] Radiation monitor vault detector zero unstable. This amplifier was the last remaining original Kiethley K302 electrometer amplifier. It has now been replaced with a Burr Brown OPA128 electrometer amplifier as the others have been replaced.